OpenStudy (anonymous):

please help Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 3, -13, and 5 + 4i

OpenStudy (anonymous):

you cannot be helped.

OpenStudy (anonymous):

only 5 + 4i? Shouldn't it be + or -?

OpenStudy (anonymous):

i can help

OpenStudy (anonymous):

A. f(x) = x4 - 8x3 - 12x2 + 400x - 1599 B. f(x) = x4 - 200x2 + 800x - 1599 B. f(x) = x4 - 98x2 + 800x - 1599 c. f(x) = x4 - 8x3 + 12x2 - 400x + 1599

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

well 5+4i is what the base equation is I'm guessing and i have to make it into a polynomial function with the 3 and -13

OpenStudy (anonymous):

Remember, if the polynomial has the given zeroes, then first start off writing it in factored form:\[f(x)=(x-3)(x+13)(x-(5+4i))...\]So we know the function will look something like that in factored form. But if you recall, then you would know that there is always conjugate complex roots of polynomials, i.e if a + bi is a root then a - bi must also be a root. Therefore, the complete factored form of the function with least degree will be:\[f(x)=(x-3)(x+13)(x-(5+4i))(x-(5-4i))\]Now just expand that function and you'll get the polynomial with least degree with the given zeroes. Can you do that? @Andresshoyos23

OpenStudy (anonymous):

@Andreshoyos23